What is the slope of the line 3x= 5y-3?

Found 2 solutions by stanbon, tutorcecilia:
Answer by stanbon(75887)

slope of the line 3x= 5y-3?
5y=3x+3
y = (3/5)x + (3/5)
------------
slope = coefficient of "x" which is (3/5)
===========================
Cheers,
Stan H.

Answer by tutorcecilia(2152)

3x= 5y-3? [solve or isolate the y-term]
3x+3=5y-3+3
3x+3=5y
3x/5+3/5=5y/5
(3/5)x+3/5=y [re-arrange the terms]
y=(3/5)x+3
The slope is the m-term in the slope-intercept form of a line: y=mx+b
Slope=m=3/5